# Preprint submitted to Computers & Geosciences

### v4

@@ 3 Geomodel generation and flow simulations using diffusion models
We now consider quantitative similarity metrics. First, we compute two-point connectivity functions for the two sets of realizations. This function gives the probability that two points separated by a given lag distance, along a specified direction, belong to the same facies (Laloy et al., 2018). Results are presented in Figure 6 for the mud, levee, and channel facies along the horizontal and diagonal ($y=x$) directions. We observe near-perfect agreement for the P10, P50, P90 (percentile) results for these quantities. Rather than show these percentile results, we follow Laloy et al. (2018) and present mean, maximum, and minimum values at each lag for the reference Petrel models, along with individual curves for LDM-generated models. The general agreement between the two sets of results in Figure 6 indicates that the overall orientation and connectivity for the mud, levee, and channel facies are accurately captured in the LDM realizations.
@@
The similarity between models can be quantified using the structural similarity index measure (SSIM), a widely applied image similarity metric (Wang et al., 2004). SSIM ranges from 1 (for identical images) to -1. Results for the two cases are shown in Figure 8. In these results we use increments in $\delta$ of 0.05. The red curves provide SSIM computed between consecutive models (i.e., separated by $\delta=0.05$), while the blue curves give SSIM between the first model $\mathbf{m}_{DM}^{(1)}$ and the model corresponding to each $\delta$ value. Although there is some variability in the red curves we do observe clear consistency from model to model, with SSIM values of around 0.8. The blue curves display a smooth transition away from $\mathbf{m}_{DM}^{(1)}$. Both sets of behaviors are desirable in history matching settings and suggest that the LDM should be applicable in this context.
@@
The setup for the flow problem is as follows. The geomodels contain $64\times 64$ cells, with each grid block of dimensions 20 m $\times$ 20 m $\times$ 5 m (in the $x$, $y$ and $z$ directions, respectively). We consider two-phase oil-water flow. The relative permeability curves used in all simulations are shown in Figure 9(a). Water viscosity is constant at 0.31 cP. Oil viscosity is a function of pressure, with a value of 1.09 cP at 325 bar. The initial water saturation and initial reservoir pressure are 0.1 and 310 bar. Porosity and permeability are constant within each facies, at 0.2 and 2500 mD for the channel facies, 0.15 and 400 mD for the levee facies, and 0.05 and 50 mD for the mud facies. There are a total of five wells – three injection wells and two production wells – at the conditioning locations, as shown in Figure 9(b). Wells are BHP controlled, with injectors at 330 bar and producers at 300 bar. The simulations are run for a period of 2500 days, with maximum time steps of 50 days. All runs are performed using Stanford’s Automatic Differentiation General Purpose Research Simulator, ADGPRS (Zhou, 2012).
@@ 4 History matching using the LDM
The history matching algorithm used in this work is the ensemble smoother with multiple data assimilation, ESMDA, developed by Emerick and Reynolds (2013). ESMDA is widely applied for history matching with high-dimensional models, i.e., in the absence of any geological parameterization. In such cases, heuristic treatments such as localization are commonly used. ESMDA has also been applied for history matching with parameterized models. Examples include Canchumuni et al. (2019, 2021), Liu and Durlofsky (2021), and Jiang and Durlofsky (2021). ESMDA is well-suited for use with the LDM parameterization because the latent variable is (close to) standard normal, consistent with the Gaussian assumptions inherent in ESMDA.
@@
where the overbar denotes the average over the $N_{e}$ samples in the ensemble at the current iteration. At each iteration, the observed data are randomly perturbed and the inflation factor is varied. Inflation factors in ESMDA must satisfy $\sum_{i=1}^{N_{a}}{\alpha_{i}}^{-1}=1$. In this work, we use an ensemble size of $N_{e}=200$, $N_{a}=10$ assimilation steps, and inflation coefficients $\alpha_{i}$, $i=1,\ldots,10$, of [57.017, 35.0, 25.0, 20.0, 18.0, 15.0, 12.0, 8.0, 5.0, 3.0], as suggested by Emerick and Reynolds (2013).
@@ 4.2 Case 2: uncertain facies properties
Facies Porosity Permeability [$\bm{\log_{e}k}$ mD]

### v5

@@ 3 Geomodel generation and flow simulations using diffusion models
We now consider quantitative similarity metrics. First, we compute two-point connectivity functions for the two sets of realizations. This function gives the probability that two points separated by a given lag distance, along a specified direction, belong to the same facies (Laloy_2018). Results are presented in Figure 6 for the mud, levee, and channel facies along the horizontal and diagonal ($y=x$) directions. We observe near-perfect agreement for the P10, P50, P90 (percentile) results for these quantities. Rather than show these percentile results, we follow Laloy_2018 and present mean, maximum, and minimum values at each lag for the reference Petrel models, along with individual curves for LDM-generated models. The general agreement between the two sets of results in Figure 6 indicates that the overall orientation and connectivity for the mud, levee, and channel facies are accurately captured in the LDM realizations.
@@
The similarity between models can be quantified using the structural similarity index measure (SSIM), a widely applied image similarity metric (ssim). SSIM ranges from 1 (for identical images) to -1. Results for the two cases are shown in Figure 8. In these results we use increments in $\delta$ of 0.05. The red curves provide SSIM computed between consecutive models (i.e., separated by $\delta=0.05$), while the blue curves give SSIM between the first model $\mathbf{m}_{DM}^{(1)}$ and the model corresponding to each $\delta$ value. Although there is some variability in the red curves we do observe clear consistency from model to model, with SSIM values of around 0.8. The blue curves display a smooth transition away from $\mathbf{m}_{DM}^{(1)}$. Both sets of behaviors are desirable in history matching settings and suggest that the LDM should be applicable in this context.
@@
The setup for the flow problem is as follows. The geomodels contain $64\times 64$ cells, with each grid block of dimensions 20 m $\times$ 20 m $\times$ 5 m (in the $x$, $y$ and $z$ directions, respectively). We consider two-phase oil-water flow. The relative permeability curves used in all simulations are shown in Figure 9(a). Water viscosity is constant at 0.31 cP. Oil viscosity is a function of pressure, with a value of 1.09 cP at 325 bar. The initial water saturation and initial reservoir pressure are 0.1 and 310 bar. Porosity and permeability are constant within each facies, at 0.2 and 2500 mD for the channel facies, 0.15 and 400 mD for the levee facies, and 0.05 and 50 mD for the mud facies. There are a total of five wells – three injection wells and two production wells – at the conditioning locations, as shown in Figure 9(b). Wells are BHP controlled, with injectors at 330 bar and producers at 300 bar. The simulations are run for a period of 2500 days, with maximum time steps of 50 days. All runs are performed using Stanford’s Automatic Differentiation General Purpose Research Simulator, ADGPRS (zhou2012parallel).
@@ 4 History matching using the LDM
The history matching algorithm used in this work is the ensemble smoother with multiple data assimilation, ESMDA, developed by EMERICK20133. ESMDA is widely applied for history matching with high-dimensional models, i.e., in the absence of any geological parameterization. In such cases, heuristic treatments such as localization are commonly used. ESMDA has also been applied for history matching with parameterized models. Examples include Canchumuni_2019; Canchumuni_2020, LIU2021104676, and sujiang_2021. ESMDA is well-suited for use with the LDM parameterization because the latent variable is (close to) standard normal, consistent with the Gaussian assumptions inherent in ESMDA.
@@
where the overbar denotes the average over the $N_{e}$ samples in the ensemble at the current iteration. At each iteration, the observed data are randomly perturbed and the inflation factor is varied. Inflation factors in ESMDA must satisfy $\sum_{i=1}^{N_{a}}{\alpha_{i}}^{-1}=1$. In this work, we use an ensemble size of $N_{e}=200$, $N_{a}=10$ assimilation steps, and inflation coefficients $\alpha_{i}$, $i=1,\ldots,10$, of [57.017, 35.0, 25.0, 20.0, 18.0, 15.0, 12.0, 8.0, 5.0, 3.0], as suggested by EMERICK20133.
@@ 4.2 Case 2: uncertain facies properties
Facies Porosity Permeability $\bm{\log_{e}k}$ mD]